How long is a guy wire reaching from the top of a 18ft pole to a point on the ground 12 ft from the pole.
What is the exact length in feet, using radicals as needed?
What is the length approximated to three decimal places? Feet, round to the nearest thousandths. Thank you so much!
: How long is a guy wire reaching from the top of a 18ft pole to a point on the ground 12 ft from the pole.
What is the exact length in feet, using radicals as needed?
What is the length approximated to three decimal places? Feet, round to the nearest thousandths. Thank you so much!

Answer by checkley77(1746)

HERE WE HAVE A RIGHT TRIANGLE WITH SIDES OF 18 & 12 FEET.
USING THE EQUATION A^2+B^2=C^2
WE HAVE:
18^2+12^2=C^2
324+144=C^2
468=C^2
C=SQRT468
C=21.6333 OR 21.633 FEET FOR THE LENGTH OF THE GUY WIRE.